OpenStudy (rainbow_dash):

Find the sum of the given infinite geometric series. 2 + 1 + 1/2 + 1/4 + ...

OpenStudy (anonymous):

3 3/4

OpenStudy (anonymous):

\[\Huge 2+1+\frac{1}{2} + \frac{1}{4}+ ... = \sum_{n=0}^{\infty}\frac{2}{2^n}= 2\sum_{n=0}^{\infty}\frac{1}{2^n}\]